OpenStudy (anonymous):

(I will give medal and fan you!~)(Want a,b,c or d please...? XD) (OCA Problem... plz help im really behind Giving Medal And Fan~!) Caroline runs 3 1/2 miles in 22 1/4 minutes. How many miles does she run per minute? A) 7/12 mile per minute B) 1 5/7 mile per minute C) 1 1/15 mile per minute D) 1 4/11 mile per minute

OpenStudy (anonymous):

lets make them improper fractions first 3 1/2=7/2 22 1/4=89/4 rate=distance/time so rate=(7/2)/(89/4) when you divide by a fraction multiply by the reciprocal so 7/2 * 4/89=?